Shift Summary for owl shift, 12/15/99

John Shields (jshields@klak)
Wed, 15 Dec 1999 07:01:22 -0600

Shift Summary for owl shift, 12/15/99

On Shift: J. Shields, S. Taegar

Hours of beam available at all intensities: 6.5
Hours of beam available at nominal intensity: 6.5
Hours of E799 data (special runs not included): 5.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

NM2Q1 trips off, brief access to reset CAEN crate, DAQ_DEAD, swapping tapes
========================================================================
Runs ended this shift:
Run 15219 - 26.8M - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift: 1

reset CAEN power supply to restore TRD monitoring.

========================================================================
Overview of shift:

E799 datataking, NM2Q1 trips off, atev3 dies, Sydney & Valeri go on a quick
access, more datataking, pipeline/trigger/DAQ_DEAD problems, tapes
changed.
Also, "scalar out of limits" problems persist throughout the shift,
probably at least in part due to the large fluctuation in SEM between
spills (varied from 8e12 - 1.1e13, though the majority were near nominal)

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15219 01:55 NM2Q1 trips, beam is lost. ESt. is 10 min
Valeri is contacted.

02:20 Valeri arrives, notices atev3 has died.
atev3 is then rebooted, and Valeri & Sydney
go on access to power cycle the CAEN crate.

02:25 NM2Q1 reset. Beam to users, but it disappears
after one spill anyway.

02:40 VJ & ST return from access.

02:46 beam returns to ktev

618 05:05 pipeline errors, init hangs on crate 1. Crate 1
rebooted, then init/kfrend/power cycle repeater/
kfrend.

05:15 triggers resume

05:16 notice that DAQMON stopped updating at spill 503.
Ran reaper & snapshot commands

05:30 DAQ_DEAD errors for most of spill. ktev 2 is
processing event late into next spill. Suspect
that tapes may be full and frozen DAQMON won't
tell us.

05:35 stop run, but run control hangs. Did l3#save
and drt_frest_start. Changed tapes to
UP*226-228, UPS076. Also did a SLOWDA fresh
start, as per Valeri's request.

06:00 start run 15220

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift